OpenStudy (anonymous):

help with logs!!! 1. e^2x-4e^x+3=0 2. 2log(base3)(x+4)=log(base3)9 + 2 3. log(base2)(x-6) + log(base2)(x-4)-log(base2)x=2 4. x^2x + 2^x - 12 = 0 5. 2 ^ 2x + 3^x - 2 = 0 6. e^2x - 2 e^x - 3 = 0

OpenStudy (anonymous):

I) put y=e^x then your eqn will be y^2-4y+3=0 use Quadratic equation to solve this

OpenStudy (anonymous):

2log(base3)(x+4)=log(base3)9 + 2 Implies 2log(base3)(x+4)-log(base3)9=2 log(base3){[(x+4)^2]/9}=2 then exponentiate wit 3 and solve

OpenStudy (anonymous):

same as above for 3

OpenStudy (anonymous):

what do you mean by exponentiate?

OpenStudy (anonymous):

if log(basek)m=n then k^n=m

OpenStudy (anonymous):

so 3^2 = (x+4)^2/9?

OpenStudy (anonymous):

for example in your qn log(base3){[(x+4)^2]/9}=2 on exponentiation [(x+4)^2]/9=3^2

OpenStudy (anonymous):

Yeah you r doing good

OpenStudy (anonymous):

i have x^2 + 8 x - 65 = 0 is that right?

OpenStudy (anonymous):

@Joseph91